# A molecular dynamics study of size effects for critical resolved shear stress in nickel superalloys

### v2

@@ Abstract
We present in this work a molecular dynamics study of a size effect relating to the volume fraction of precipitate of edge dislocation motion in a simple model of Nickel superalloys. We model the superalloy as periodically spaced cubic precipitates inside a uniform matrix. We then analyze the motion of paired edge dislocations in the phase when subject to an external shear stress for various volume fractions of the precipitate for a wide range of temperatures, from 300 K to 700 K. While the variation of dislocation velocity is not significant, the critical resolved shear stress is found to exhibit a power law dependence on the volume fraction of the precipitate with two distinct regimes which have similar exponent but markedly different prefactors; we also observe that this two-regime behavior remains true across a wide range of temperatures. We present a detailed analysis of this behavior and reduce it to a linear dependence of the critical resolved shear stress on the length of the precipitate along the direction of dislocation motion. We further identify the critical length scale underlying the transition between the two observed regimes as the total core width of the paired dislocations in a pure system, which includes in addition to the complex stacking fault separating the partials of the paired dislocations the width of the anti-phase boundary that is formed between the super-dislocations. The results presented in this work provides new details on the strengthening effect of precipitates in nickel superalloys and also has important implications for larger scale dislocation dynamics studies for nickel superalloys.
[IITB]organization=Department of Aerospace Engineering, Indian Institute of Techhnology Bombay, city=Mumbai, state=Maharashtra 400076, country=India
[Cal Poly]organization=Aerospace Engineering, California Polytechnic State University, city=San Luis Obispo, state=CA 93407, country=USA
@@ 1 Introduction
In the past decade, numerous experimental studies have explored how microstructural features affect the mechanical properties of these alloys [29, 9]. Transmission electron microscopy investigations indicate clear evidence of dislocation interactions with precipitates, revealing that dislocation lines tend to bend at the interface [26, 6]. The study by [18] showed that the variation in lattice parameters between the and phases increases with temperature, particularly below 1100∘C, thereby significantly impacting dislocation mobility. Dislocation motion is further hindered by interfacial dislocation networks, which enhance creep resistance [5, 38]. Several mechanisms contribute to precipitate strengthening, such as (1) chemical strengthening that arises from increased surface energy resistance during dislocation shearing, (2) stacking fault strengthening due to variations in stacking fault energy, (3) modulus strengthening resulting from mismatches in elastic moduli, (4) coherency strain strengthening stemming from differences in lattice parameters, and (5) order strengthening linked to disruptions in the ordered Ni3Al crystal structure during dislocation motion [3]. Although, [28] suggests that the role of the coherency strain effect is negligible, whereas [21] suggests that it plays a significant role. Among these strengthening mechanisms, the interplay between dislocation shearing and bypassing appears to be crucial. The extent research on this topic indicates that the size of the precipitate is a decisive factor in determining whether dislocations will shear or bypass them [11, 16, 1]. Dislocations generally shear smaller precipitates, with a mean radius smaller than 100–120 Å, and bypass larger ones via the Orowan mechanism [32, 10]. In -hardened superalloys such as NIMONIC PE16, the primary deformation mechanism of dislocations interacting with precipitates evolves during aging. For NIMONIC PE16 with a volume fraction , the transition occurs at an average particle radius of approximately –, corresponding to a critical resolved shear stress –. Below this radius, dislocation shearing dominates, leading to an increase in strength; above it, bypassing via Orowan looping becomes favorable, resulting in a decrease in [22]. The critical precipitate radius is often utilized as a crucial factor to indicate the transition between these two mechanisms [20, 10, 19, 7]. However, experimental methods alone are inadequate to resolve all the details of the atomistic interactions, highlighting the need for MD simulations.
@@
Several MD studies have investigated the nature of dislocation-precipitate interactions in Nickel superalloys. Early MD simulations indicated that dislocations undergo a transition from shearing to Orowan looping as the size of the precipitates increases [24]. Studies by [13, 15] confirmed this transition. The coupling is typically classified as strong or weak depending on the size and distribution of the precipitates—these factors directly influence the CRSS for dislocation motion. A key finding from MD studies is that the morphology and distribution of precipitates can lead to an intermediate strengthening regime where dislocation behavior cannot be clearly categorized into either shearing or bypassing [37]. This aspect is particularly pertinent for ellipsoidal precipitates in IN718, where existing strengthening models fail to capture the morphology-dependent dislocation interactions [36]. Bacon and Osetsky [4] explored the interactions between edge dislocations and copper precipitates in -iron alloys, demonstrating how the CRSS depends on factors such as temperature, precipitate size, and lattice mismatch. Similar results have been documented for medium- and high-entropy alloys [23, 2, 17], as well as Mg-Al alloys [34] and Mg-Zn alloys [8], showing that smaller precipitates tend to promote dislocation shearing, whereas larger ones promote the bypassing mechanism. Beyond precipitate interactions, the intrinsic mobility of dislocations under local stress states plays a critical role in plasticity. Recent MD studies by [30] quantified how non-Schmid stresses alter dislocation velocities in aluminum, revealing character-angle-dependent damping coefficients and nonlinear mobility regimes.
@@
While prior experimental and computational studies have explored dislocation-precipitate interactions, we present new scaling behavior for dislocation-precipitate interactions in Nickel superalloys based on MD simulations that have not been reported earlier, to the best of our knowledge. We present a detailed MD study of the interaction of paired edge dislocations in a system with cuboidal precipitates, focusing on the effects of volume fraction (1–60%) and temperature (300–700 K). We demonstrate that CRSS follows a power-law scaling with volume fraction, exhibiting two distinct regimes separated by a critical precipitate size . This length scale corresponds to the total core width of paired dislocations in the phase (encompassing complex stacking faults(CSF) and anti-phase boundaries(APB)) and remains independent of temperature over the temperature range indicated above. We also establish that this behaviour reflects a linear relationship between CRSS and the precipitate length along the direction of dislocation motion. These findings thus provide new size effects for dislocation-precipitate interactions in Ni-based superalloys that are useful for mesoscale dislocation dynamics models and for alloy design studies.
@@ 2.1 Dislocation Mobility in Ni and Ni3Al
The dislocation velocity as a function of the applied stress is shown in Figure 2(a) for the phase, and in Figure 2(b) for the phase. A comparison of the dislocation velocities in the and phases as a function of the applied stress is shown in Figure 3. The details of the computational setup used to obtain these results are summarized in A. For stress values under about 100 MPa, the dislocation velocity increases steeply with the applied stress in both the and phases. For higher values of stress, the dislocation velocities saturate, with the velocity in the phase being larger than that in phase; this is consistent with the higher energy barriers in the ordered structure of the phase. As can be seen from Figure 3, the difference in the saturation velocities in the and phases are not significant. Thus, to a first degree of approximation, the dislocation velocity can be approximated as being equal in both the phases.
@@ 2.2 Dislocation Motion in System
The setup of the simulation used to analyze the interaction of edge dislocations with periodically placed precipitates in a matrix is illustrated in Figure 4. To simulate the experimental observation that dislocations travel in pairs in + systems, we initialized two edge dislocations separated by distance in the matrix. We refer to the leading and trailing dislocations in the dislocation pair as and dislocations, respectively. The precipitates are modeled as cuboidal regions of side length that are periodically distributed with a center-to-center spacing of along the direction of dislocation motion, which we refer to as the direction. We varied the volume fraction of , , from to to understand the interaction of the paired dislocations in the phase with the precipitate. The dislocation velocity in the + system as a function of the applied stress for various volume fractions of the precipitate is shown in Figure 5. The velocities of both the and dislocations are indicated in Figure 5—these are nearly identical and thus can be taken to be velocity of the paired dislocations. It is also evident from Figure 5 that while the saturation velocity for the paired dislocations is similar across different volume fractions of the precipitate, the CRSS values vary significantly.
@@ 4 Conclusion
In this work, we performed extensive molecular dynamics simulations of dislocation motion in Nickel superalloys, focusing in particular on the effect of the size of the precipitate. Simulations involving the motion of paired dislocations in phases containing cuboidal precipitates revealed several interesting observations: (i) the dislocation velocity does not vary significantly as a function of precipitate size, (ii) the critical resolved shear stress exhibits a power law relationship with the volume fraction, with two clearly demarcated regimes, and (iii) a further investigation of these trends indicate that the dependence of CRSS is in fact linear as a function of the side length of the precipitate. Further, the two distinct regimes are separated by a critical length scale Å, which correlates with the total width of the paired dislocation core in the pure phase. We presented several snapshots of paired dislocation motion in systems to highlight this behavior. We further presented qualitative differences in the separation of the paired dislocations in the two regimes. These results have important implications for dislocation dynamics simulations, and will be pursued in a future work.
Appendix A Methods
@@
We used the LAMMPS code [33] to perform all the molecular dynamics simulations, Atomsk [12] to generate various atomic configurations, and Ovito [31] for visualizing atomic configurations. For all the simulations, we use the embedded atom interatomic potential developed in [25]. We use the model developed in [23] to simulate straight edge dislocations. The following coordinate directions are chosen: for the X-axis, for the Y-axis, and for the Z-axis. The simulated model contains close to 909,696 atoms, with dimensions of along the X, Y, and Z axes, respectively. Periodic boundary conditions are used along the X- and Z-axes. To compute the dislocation velocities, we equilibrate the initial configuration and follow the procedure given in [14]. For minimizing energy of the initial configurations, we use the conjugate gradient algorithm; the energy minimization process eliminates any high-energy configurations and ensures that the initial configuration is stable and ready for dynamic simulations. In all simulations, we maintain a constant temperature in the model using a Nose-Hoover thermostat. All simulations were equilibrated at the desired fixed temperature for 100,000-time steps, with a timestep of 0.001. The simulation box comprises of three distinct sections as shown in Figure 4: the rigid region (R), the thermostatted region () at the top and bottom, and the central mobile area (M). The rigid region (R) covers of the total length () along the Y-axis, denoted by = —all atoms were modeled as individual rigid bodies in this region. The central part of our simulation box, designated as the mobile region (M), spans of . We employed an NVE thermostat for both the rigid and mobile regions to conserve energy during the simulations. Adjacent to the mobile region, on both the top and bottom are the thermostat regions (), with . These thermostat regions are crucial for maintaining thermal equilibrium in the simulation. In these regions, the NVT ensemble was applied to maintain the system at the desired constant temperature, and further compensates for the heat that is generated due to dislocation motion during shearing. Shear stresses are applied across the rigid regions in the top and bottom of the simulation box. For the shear force, we apply a force using the equation: , where and are dimensions of the simulation box along respective directions, and is the number of atoms in the corresponding rigid region. For the top rigid region, the force is applied in the positive -direction and in the negative -direction for the bottom rigid region to create the necessary shear stresses for our simulations.

### v3

# Size-Dependent Power Laws for Edge Dislocations in Nickel Superalloys: A Molecular Dynamics Study

@@ Abstract
We present in this work computational evidence, using molecular dynamics simulations, of a size effect in the relationship between the critical resolved shear stress for edge dislocation motion in nickel superalloys and the size of precipitates, under certain conditions. We model the superalloy as periodically spaced cubic precipitates inside a uniform matrix. We then analyze the motion of paired edge dislocations in the phase when subject to an external shear stress for various volume fractions of the precipitate for a wide range of temperatures, from 300 K to 700 K. While the variation of dislocation velocity is not significant, the critical resolved shear stress is found to exhibit a power law dependence on the volume fraction of the precipitate with two distinct regimes which have similar exponent but markedly different prefactors; we also observe that this two-regime behavior remains true across a wide range of temperatures. We present a detailed analysis of this behavior and reduce it to a linear dependence of the critical resolved shear stress on the length of the precipitate along the direction of dislocation motion. We further identify the critical length scale underlying the transition between the two observed regimes as the total core width of the paired dislocations in a pure system, which includes in addition to the complex stacking fault separating the partials of the paired dislocations the width of the anti-phase boundary that is formed between the super-dislocations. We present auxillary results using spherical precipitates that exhibit the same trend, but a full analysis of the interplay between size of the precipitate, volume fraction, and other related factors is not pursued in this work. Despite the special configurations considered in this work, the results presented here highlights non-trivial size-dependent effects, provides new details on the strengthening effect of precipitates in nickel superalloys, and has important implications for larger scale dislocation dynamics studies for nickel superalloys.
@@ 1 Introduction
In the past decade, numerous experimental studies have explored how microstructural features affect the mechanical properties of these alloys [29, 9]. Transmission electron microscopy investigations indicate clear evidence of dislocation interactions with precipitates, revealing that dislocation lines tend to bend at the interface [26, 6]. The study by [18] showed that the variation in lattice parameters between the and phases increases with temperature, particularly below 1100∘C, thereby significantly impacting dislocation mobility. Dislocation motion is further hindered by interfacial dislocation networks, which enhance creep resistance [5, 38]. Several mechanisms contribute to precipitate strengthening, such as (1) chemical strengthening that arises from increased surface energy resistance during dislocation shearing, (2) stacking fault strengthening due to variations in stacking fault energy, (3) modulus strengthening resulting from mismatches in elastic moduli, (4) coherency strain strengthening stemming from differences in lattice parameters, and (5) order strengthening linked to disruptions in the ordered Ni3Al crystal structure during dislocation motion [3]. However, Raynor et al. [28] suggest that the role of the coherency strain effect is negligible, whereas Miller et al. [21] suggests that it plays a significant role. Among these strengthening mechanisms, the interplay between dislocation shearing and bypassing appears to be crucial. The extent research on this topic indicates that the size of the precipitate is a decisive factor in determining whether dislocations will shear or bypass them [11, 16, 1]. Dislocations generally shear smaller precipitates, with a mean radius smaller than 100–120 Å, and bypass larger ones via the Orowan mechanism [32, 10]. In -hardened superalloys such as NIMONIC PE16, the primary deformation mechanism of dislocations interacting with precipitates evolves during aging. For NIMONIC PE16 with carbon content , the transition occurs at an average particle radius of approximately –, corresponding to a critical resolved shear stress –. Below this radius, dislocation shearing dominates, leading to an increase in strength; above it, bypassing via Orowan looping becomes favorable, resulting in a decrease in [22]. The critical precipitate radius is often utilized as a crucial factor to indicate the transition between these two mechanisms [20, 10, 19, 7]. However, experimental methods alone are inadequate to resolve all the details of the atomistic interactions, highlighting the need for MD simulations.
@@
Several MD studies have investigated the nature of dislocation-precipitate interactions in Nickel superalloys. Early MD simulations indicated that dislocations undergo a transition from shearing to Orowan looping as the size of the precipitates increases [24]. Studies by Hocker et al. [13] and Kohler et al. [15] confirmed this transition. The coupling is typically classified as strong or weak depending on the size and distribution of the precipitates—these factors directly influence the CRSS for dislocation motion. A key finding from MD studies is that the morphology and distribution of precipitates can lead to an intermediate strengthening regime where dislocation behavior cannot be clearly categorized into either shearing or bypassing [37]. This aspect is particularly pertinent for ellipsoidal precipitates in IN718, where existing strengthening models fail to capture the morphology-dependent dislocation interactions [36]. Bacon and Osetsky [4] explored the interactions between edge dislocations and copper precipitates in -iron alloys, demonstrating how the CRSS depends on factors such as temperature, precipitate size, and lattice mismatch. Similar results have been documented for medium- and high-entropy alloys [23, 2, 17], as well as Mg-Al alloys [34] and Mg-Zn alloys [8], showing that smaller precipitates tend to promote dislocation shearing, whereas larger ones promote the bypassing mechanism. Beyond precipitate interactions, the intrinsic mobility of dislocations under local stress states plays a critical role in plasticity. Recent MD studies by Shen et al. [30] quantified how non-Schmid stresses alter dislocation velocities in aluminum, revealing character-angle-dependent damping coefficients and nonlinear mobility regimes.
@@
While prior experimental and computational studies have explored dislocation-precipitate interactions, we present new scaling behavior for dislocation-precipitate interactions in Nickel superalloys based on MD simulations that have not been reported earlier, to the best of our knowledge. We present a detailed MD study of the interaction of paired edge dislocations in a system with cuboidal precipitates, focusing on the effects of volume fraction (1–60%) and temperature (300–700 K). We demonstrate that CRSS follows a power-law scaling with volume fraction, exhibiting two distinct regimes separated by a critical precipitate size . This length scale corresponds to the total core width of paired dislocations in the phase (encompassing complex stacking faults(CSF) and anti-phase boundaries(APB)) and remains independent of temperature over the temperature range indicated above. We also establish that this behaviour reflects a linear relationship between CRSS and the precipitate length along the direction of dislocation motion. We also present auxillary simulations with spherical precipitates that reveal the same trend. These findings thus provide new size effects for dislocation-precipitate interactions in Ni-based superalloys that are useful for mesoscale dislocation dynamics models and for alloy design studies.
@@ 2.1 Dislocation Mobility in Ni and Ni3Al
The dislocation velocity as a function of the applied stress is shown in Figure 2(a) for the phase, and in Figure 2(b) for the phase. A comparison of the dislocation velocities in the and phases as a function of the applied stress is shown in Figure 3. The details of the computational setup used to obtain these results are summarized in Appendix A. For stress values under about 100 MPa, the dislocation velocity increases steeply with the applied stress in both the and phases. For higher values of stress, the dislocation velocities saturate, with the velocity in the phase being larger than that in phase; this is consistent with the higher energy barriers in the ordered structure of the phase. As can be seen from Figure 3, the difference in the saturation velocities in the and phases are not significant. Thus, to a first degree of approximation, the dislocation velocity can be approximated as being equal in both the phases. These results indicate that since the dislocation velocity is roughly constant, the more interesting quantity to analyze is the critical resolved shear stress for dislocation motion, which we turn to next.
@@ 2.2 Dislocation Motion in System
The setup of the simulation used to analyze the interaction of edge dislocations with periodically placed precipitates in a matrix is illustrated in Figure 4. To simulate the experimental observation that dislocations travel in pairs in + systems, we initialized two edge dislocations separated by distance in the matrix. We refer to the leading and trailing dislocations in the dislocation pair as and dislocations, respectively. The precipitates are modeled as cuboidal regions of side length that are periodically distributed with a center-to-center spacing of along the direction of dislocation motion, which we refer to as the X direction. We varied the volume fraction of , , from to to understand the interaction of the paired dislocations in the phase with the precipitate. The dislocation velocity in the + system as a function of the applied stress for various volume fractions of the precipitate is shown in Figure 5. The velocities of both the and dislocations are indicated in Figure 5—these are nearly identical and thus can be taken to be velocity of the paired dislocations. It is also evident from Figure 5 that while the saturation velocity for the paired dislocations is similar across different volume fractions of the precipitate, the CRSS values vary significantly.
@@ 3 Discussion
We discuss various aspects of our results and their shortcomings in this section.
3.1 Time snapshots of dislocation motion
@@ 3.2 Variations in dislocation core width
3.3 Spherical precipitates
The simulations presented so far dealt with cubic precipitates. To check if the conclusions drawn here are valid for other precipitate shapes, we simulated spherical precipitates with a diameter similar to the side length of the cubic precipitates in the original study, as shown in Figures 22 and 22. The CRSS values for edge dislocations cutting at the equatorial plane of the sphere are shown in Figure 23, and clearly indicate a jump in CRSS when the diameter of the sphere is equal to the critical length identified in the case of cubic precipitates. We performed a power-law fit, which yielded an exponent value of for , and for . Note that these exponents are close to , thereby indicating that a roughly linear relationship exists between the CRSS values and the value observed for cubic precipitates. This result further supports our conclusion that the CRSS depends on the precipitate length () and not just its volume fraction, reinforcing the robustness of our proposed relationship.
A comparison of the CRSS values for cubic and spherical precipitates is shown in Figure 24. As can be seen in this figure, the CRSS clearly transitions at the critical length scale ( Å), which corresponds to the total dislocation core width (2 CSFs + APB). We also observe that CRSS shows a linear dependence on precipitate length () for both these geometries. These simulations with different precipitate shapes provide further evidence that the CRSS is governed by the precipitate’s length, not just its volume fraction.
The results presented here do not guarantee that these trends generalize to all precipitate shapes, but they are rather presented as preliminary evidence that the two-phase power law behavior for CRSS may be applicable to a larger class of precipitates than the ones considered in this work.
3.4 Non-zero relative orientations of the and phases
An important limitation of the present work is that all the simulations have precipitates with orientations identical to that of the phase. A full parametric study involving different relative orientations between the dislocation slip plane and the precipitate is beyond the scope of this work. We studied a special case where an initially conforming spherical precipitate is rotated by about the Z-axis. We observed a significant increase in the Critical Resolved Shear Stress (CRSS) due to this rotation. For example, for a spherical precipitate with a diameter of 100 Å, the CRSS was 150 MPa when the precipitate was unrotated. However, when the precipitate was rotated, the CRSS increased to 280 MPa. This is expected since the mechanism of dislocation motion changes in this case. A full investigation of the effect of different relative orientations and the corresponding mechanisms will be pursued in a future work. It is also unclear if the same two-phase behavior observed here will carry over in such cases.
We reiterate that the intention behind this work is not to perform a full parametric sweep of all the design variables. This is both computationally expensive and outside the scope of this work. Rather, our intent here is to highlight a previously unnoticed phase transition in the CRSS behavior under certain special conditions.
3.5 Interplay of volume fraction and precipitate size
An important limitation of this work is the undersampling of the volume fraction vs size of the precipitate in the direction of dislocation motion. We have deliberately restricted the scope of our simulations to emphasize the latter. To motivate this restriction, we conducted two simulations with varying (cubic) precipitate sizes—quantified by the dimensions —but with roughly similar volume fraction . The cell size in these simulations are indicated as . The simulation results are shown in Table 3. These inform us that samples with different precipitate sizes corresponding to roughly similar volume fractions have similar CRSS values. We note that this is a small sample set to draw definitive conclusions, but since the mechanism we postulate for the size effect is independent of the dimension of the precipitate in a direction normal to the slip plane, we have not pursued more simulations to reduce the computational cost. A full investigation of the effect of volume fraction will be pursued in a future work.
CRSS | |||||||
250 | 130 | 200 | 130 | 200 | 130 | 22.21 | 280 |
130 | 60 | 100 | 60 | 60 | 50 | 23.08 | 270 |
3.6 Weak coupling limit
All the simulations presented here involved paired dislocations to accord with experimental observations. We also performed simulations with a setup identical to that presented in the manuscript but with only one dislocation is initialized in the gamma matrix, as shown schematically in Figure 25 to illustrate the importance of choosing paired dislocations. Physically, this corresponds to the weak coupling regime.
Snapshots of the motion of the edge dislocation through the phase for different stress levels and different volume fractions are shown in Figures 28(a)–29(b) in Appendix B. The first pass of the dislocation through the precipitate creates an APB which is then relieved when the periodic image dislocation passes through the precipitate again. The corresponding CRSS plots are shown in Figure 26. It is evident from Figure 26(a) that the values of the CRSS obtained are much higher than what is observed in experiments. This is expected since it is well known that dislocations prefer to travel in pairs in Nickel superalloys to overcome the high APB energy barrier. The primary conclusions of our work are thus based on the motion of paired dislocations. We note that even in the case where paired dislocations are introduced in the gamma phase, the leading dislocation still experiences an APB resistance. This is clearly borne out in our simulation results. Further, the CRSS values obtained with the paired dislocation simulations are in better agreement with the experimental data.
It is interesting to note that the dependence of the CRSS on the volume fraction follows a power law for the motion of single dislocations, as shown in Figure 27, but this does not exhibit the two-phase behavior.
@@ 4 Conclusion
In this work, we performed extensive molecular dynamics simulations of dislocation motion in Nickel superalloys, focusing in particular on the effect of the size of the precipitate. Simulations involving the motion of paired dislocations in phases containing cuboidal precipitates revealed several interesting observations: (i) the dislocation velocity does not vary significantly as a function of precipitate size, (ii) the critical resolved shear stress exhibits a power law relationship with the volume fraction, with two clearly demarcated regimes, and (iii) a further investigation of these trends indicate that the dependence of CRSS is in fact linear as a function of the side length of the precipitate. Further, the two distinct regimes are separated by a critical length scale Å, which correlates with the total width of the paired dislocation core in the pure phase. We presented several snapshots of paired dislocation motion in systems to highlight this behavior. We further presented qualitative differences in the separation of the paired dislocations in the two regimes. Preliminary simulations using spherical precipitates indicate that these conclusions remain valid for such precipitates too, but whether this remains true for arbitrary precipitate shapes will be investigated in the future. We have also chosen a specific subset of the volume fraction vs precipitate size choices to illustrate the fact that non-trivial two-phase behavior of CRSS for special configurations. A more detailed analysis of the interplay between the volume fraction and precipitate size, and also the relative orientation between the and phases are left for future work. Despite the special configurations considered in this work, the results presented here reveal interesting CRSS behavior that have important implications for dislocation dynamics simulations.
Data Availability
All simulation data can be obtained from the authors upon reasonable request.
Acknowledgment
We acknowledge the support and resources provided by the Aerospace Computing Environment (ACE) platform, developed by the Department of Aerospace Engineering, IIT Bombay, which were essential to the successful completion of this work. We also thank the anonymous reviewer for their helpful comments.
Appendix Appendix A Methods
@@ Acknowledgment
We used the LAMMPS code [33] to perform all the molecular dynamics simulations, Atomsk [12] to generate various atomic configurations, and Ovito [31] for visualizing atomic configurations. For all the simulations, we use the embedded atom interatomic potential developed in [25]. We use the model developed in [23] to simulate straight edge dislocations. The following coordinate directions are chosen: for the X-axis, for the Y-axis, and for the Z-axis. The simulated model contains close to 909,696 atoms, with dimensions of along the X, Y, and Z axes, respectively. Periodic boundary conditions are used along the X- and Z-axes. To compute the dislocation velocities, we equilibrate the initial configuration and follow the procedure given in [14]. For minimizing energy of the initial configurations, we use the conjugate gradient algorithm; the energy minimization process eliminates any high-energy configurations and ensures that the initial configuration is stable and ready for dynamic simulations. In all simulations, we maintain a constant temperature in the model using a Nose-Hoover thermostat. All simulations were equilibrated at the desired fixed temperature for 100,000-time steps, with a timestep of 0.001. The simulation box comprises of three distinct sections as shown in Figure 4: the rigid region (R), the thermostatted region () at the top and bottom, and the central mobile area (M). The rigid region (R) covers of the total length () along the Y-axis, denoted by = —all atoms were modeled as individual rigid bodies in this region. The central part of our simulation box, designated as the mobile region (M), spans of . We employed an NVE thermostat for both the rigid and mobile regions to conserve energy during the simulations. Adjacent to the mobile region, on both the top and bottom are the thermostat regions (), with . These thermostat regions are crucial for maintaining thermal equilibrium in the simulation. In these regions, the NVT ensemble was applied to maintain the system at the desired constant temperature, and further compensates for the heat that is generated due to dislocation motion during shearing. Shear stresses are applied across the rigid regions in the top and bottom of the simulation box. For the shear force, we apply a force using the equation: , where and are dimensions of the simulation box along respective directions, and is the number of atoms in the corresponding rigid region. For the top rigid region, the force is applied in the positive X-direction and in the negative X-direction for the bottom rigid region to create the necessary shear stresses for our simulations. For computing the velocity of edge dislocations in the pure phase, we used a single dislocation. For the rest of the simulations, namely the calculation of dislocation velocity in the pure phase and for the calculation of CRSS values in systems, we used paired dislocations. The CRSS value is estimated by increasing the shear stress by small increments of 10 MPa until the dislocation moves past the precipitate. The error bars indicated in the graphs are thus of uniform width of 10 MPa to reflect this.
Appendix Appendix B Auxillary results
Time snapshots of the motion of single dislocations in a system are presented here for the sake of completeness. We emphasize that this weak coupling regime is not observed in practice on account of the large CRSS values for dislocation motion.